### v2

# Prague, Czech

Abstract
We propose a reinforcement learning (RL) framework for adaptive precision tuning of linear solvers, and can be extended to general algorithms. The framework is formulated as a contextual bandit problem and solved using incremental action-value estimation with a discretized state space to select optimal precision configurations for computational steps, balancing precision and computational efficiency. To verify its effectiveness, we apply the framework to iterative refinement for solving linear systems $Ax=b$. In this application, our approach dynamically chooses precisions based on calculated features from the system. In detail, a Q-table maps discretized features (e.g., approximate condition number and matrix norm)to actions (chosen precision configurations for specific steps), optimized via an $\epsilon$-greedy strategy to maximize a multi-objective reward balancing accuracy and computational cost. Empirical results demonstrate effective precision selection, reducing computational cost while maintaining accuracy comparable to double-precision baselines. The framework generalizes to diverse out-of-sample data and offers insight into utilizing RL precision selection for other numerical algorithms, advancing mixed-precision numerical methods in scientific computing. To the best of our knowledge, this is the first work on precision autotuning with RL and verified on unseen datasets.
Keywords inference reduced-precision configurations, reinforcement learning, precision tuning, linear solvers
1 Introduction
Floating-point arithmetic (8766229, ) is widely used in machine learning and scientific computing. With the rapid development of large language models (DBLP:journals/corr/abs-2005-14165, ; hu2022lora, ; touvron2023, ) as well as large-scale GPU computing (DBLP:journals/corr/abs-1710-03740, ; peng2023fp8lmtrainingfp8large, ), numerical calculations and algorithm deployment with extensive use of floating-point arithmetic are often accompanied by a substantial increase in energy use. Therefore, using reduced-precision arithmetic in algorithms is typically required in many intensive computing kernels due to its lower communication, faster arithmetic, and energy efficiency. Seven commonly-used floating-point formats are referenced in Table 1. The NVIDIA A100 achieves up to 312 TFLOPS with half precision (FP16/BF16), 156 TFLOPS with TF32, and 9.7 TFLOPS with FP64, yielding an approximate $32\times$ speedup for low-precision formats. Further, the NVIDIA H100 (NVIDIA’s Hopper architecture) achieves up to 624 TFLOPS with fp8, offering an approximate $64\times$ speedup versus FP64 (nvidia_a100, ; nvidia_fp8, ). The work nvidia_fp8 demonstrates that the FP8 formats E4M3 and E5M2 can in some cases match FP16 accuracy with reduced compute and memory costs.
However, reduced-precision arithmetic often brings increased rounding errors and numerical instability, which can compromise the results of computations, especially at large scale (10.1177/1094342016652462, ). The idea behind mixed-precision algorithms is to design an algorithm that combines low and high precision arithmetic while retaining an acceptable level of accuracy, which remains a major obstacle in modern approximate computing paradigms (benkhalifa:hal-03978176, ; 7348659, ). Accordingly, many theoretical analyses on rounding error analysis and mixed-precision algorithms (doi:10.1137/17M1122918, ; doi:10.1137/17M1140819, ; doi:10.1137/20M1316822, ; doi:10.1137/17M1122918, ; abhl24, ) and precision tuning tools (6877460, ; 10.1145/3213846.3213862, ) have been developed, which aim to achieve a balance between precision loss and arithmetic speedup.
Theoretical analyses of the use of multiprecision arithmetic and mixed-precision algorithms have been extensively conducted in the context of iterative refinement (doi:10.1137/20M1316822, ; abhl24, ), least squares problems (doi:10.1137/19M1298263, ; doi:10.1137/20M1316822, ), and GMRES (jang:hal-05163845, ). For example, iterative refinement can be executed in three precisions, but the convergence is only theoretically guaranteed for a limited range of problems, which can be overly restrictive in practice oktay2022multistage . As such, one can instead use to precision tuning tools to find a suitable mixed-precision approach. Existing tools typically require running the program multiple times via trial-and-error, and the scalability and general ability are not verified, so they are either limited to small or medium-scale programs or suffer from efficiency issues. An ideal case for a precision tuning tool or method is to fit the model in a limited number of training sets and learn the key features that can be used to infer mixed-precision configurations for new unseen data. These challenges need to be addressed to reduce the dependency on programmer efforts.
Table 1: Key parameters of seven floating-point formats; $u$ denotes the unit-roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, and $e_{\max}$ denotes exponent of $x_{\max}$.
   $u$   $x_{\min}$   $x_{\max}$ $t$ $e_{\min}$ $e_{\max}$
bfloat16 (BF16) $3.91\times 10^{-3}$ $1.18\times 10^{-38}$ $3.39\times 10^{38}$ 8 -126 127
half precision (FP16) $4.88\times 10^{-4}$ $6.10\times 10^{-5}$ $6.55\times 10^{4}$ 11 -14 15
TensorFloat-32 (TF32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $70\times 10^{38} 11 -$126 127
single (FP32) $5.96\times 10^{-8}$ $1.18\times 10^{-38}$ $3.40\times 10^{38}$ 24 -126 127
double (FP64) $1.11\times 10^{-16}$ $2.23\times 10^{-308}$ $1.80\times 10^{308}$ 53 -1022 1023
A contextual bandit (also known as a multi-armed bandit with context) is a simplified RL problem where the agent makes decisions based on the current context (i.e., state) without considering the future contexts or rewards of its actions; in other words, the agent optimizes for one-step (immediate) reward rather than long-term cumulative reward. In this paper, we propose a general framework for applying reinforcement learning (RL) for precision selection in general numerical algorithms, formulated as a contextual bandit problem due to the independence of problem instances and the single decision per instance. This is the first work to formally define the use of RL for precision selection in any algorithm. The framework uses an epsilon-greedy strategy, controlled by a decaying epsilon determined by training episodes, to select precision levels for the prespecified algorithm steps, treating precision choices as a multi-armed bandit problem within a RL environment. The epsilon parameter balances exploration and exploitation, starting high to encourage diverse precision selections and decaying over time to favor optimal choices. We believe this framework enables at least two utility points:
  1. 1.
    Identifying the precision level required for user-specific computational steps in terms of hand-crafted features or low computational complexity features modeled from linear systems;
  2. 2.
    Finding salient features of the linear systems (e.g., condition number, matrix norm, non-zero values, diagonal dominance) to determine the reduced-precision configurations;
  3. 3.
    Being able to inference reduced mixed-precision configurations of unseen datasets.
We apply this framework to the GMRES-based iterative refinement (GMRES-IR) method doi:10.1137/17M1122918 to verify the effectiveness and generalization ability of RL for precision selection, aiming to achieve a trade-off between the use of low precision and accuracy. The framework is particularly suited for scenarios with limited training data and is designed to generalize to new data, and can be easily implemented in an online learning routine to avoid model retraining.
The rest of the paper is structured as follows. Section 2 discusses the existing literature on precision autotuning as well as its pros and cons, and a comparison with our work. Section 3 presents our general framework for precision selection; To verify its effectiveness, we use the iterative refinement method, a well-studied mixed-precision method, as a case study detailed in Section 4. Section 5 presents the experiments and empirical results on both dense and sparse matrices under varying condition numbers and sparsity. Section 7 discusses limitations as well as future work.
2 Related Work
There exists a variety of theoretical rounding error analyses of basic mixed-precision algorithms aimed at understanding the underlying arithmetic behavior; see the survey doi:10.1177/10943420211003313 and references therein. However, for real-world problems, which can be more complex, the scientific community typically uses more general approaches such as precision tuning tools, which operate at program compile time and runtime, immediately resulting in reduced-precision programs produced automatically. In this section, we briefly review existing approaches for the analysis of mixed-precision algorithms and automated precision tuning that are closely related to our method.
Theoretical analysis aims to design reduced-precision settings for specific steps and analyze their rounding error so that the reduced-precision arithmetic can be used safely under certain assumptions to retain algorithm accuracy and numerical stability. Analyses in WILKINSON1966 and 10.1145/321386.321394 demonstrate that in standard iterative refinement, residuals computed in extra precision (typically twice the working precision) can guarantee the convergence of the relative error to the level of the working precision if the matrix $A$ has a condition number safely less than the reciprocal of the unit-roundoff, $u$. More recently, doi:10.1137/17M1122918 found that it is possible to produce solutions to iterative refinement with normwise relative error of order $u$ for systems with condition numbers of order $u^{-1}$ or larger by using a GMRES-based approach called GMRES-IR (see Algorithm 2). The authors extended this work in doi:10.1137/17M1140819 to develop a general analysis of iterative refinement using up to three difference precisions, as well as a general solver. This work shows that, depending on constraints on the condition number $\kappa(A)$, the most expensive part of the computation can often be carried out entirely in low precision without harming the accuracy. However, to obtain theoretical results such as these, a great amount of effort is needed, and the analysis often relies on knowledge about the problem, e.g., the condition number, which may not be readily available in practice. To circumvent these obstacles, one can switch to precision tuning tools that operate during compile time and runtime; see the survey 10.1145/3381039 and references therein.
Another approach is to use an automated precision tuning method. Such methods, built upon the software level via static and dynamic analysis tools (benkhalifa:hal-03978176, ), can obtain suitable reduced-precision settings through trial and error. The search-based tool Precimonious (6877460, ) uses high-performance dynamic programming approaches, which achieve a potentially lower precision setting for the program via Delta Debugging of the space of possible floating-point configurations within the target accuracy. Further, HiFPTuner (10.1145/3213846.3213862, ) (High-level Floating-Point Tuner) operates at the source level to automatically tune the floating-point precision via profiling floating-point operations and estimating the numerical impact of the reduced-precision program. CRAFT (10.1177/1094342016652462, ) builds reduced-precision configurations of floating-point programs based on the Dyninst binary instrumentation framework (10.1177/109434200001400404, ), and uses an automatic search process that identifies the precision-level sensitivity of various parts of an application. Search-based tools for finding a low-precision configuration typically require a dramatic increase in computational cost as the number of variables increases, and do not scale well to large programs. To further reduce the runtime of precision tuning, Blame Analysis (Rubio2016, ) is devoted to executing a program once, analyzing floating-point instructions to determine the minimum precision required for each operand, reducing the search space of Precimonious and improving scalability. PyFloT (9355251, ) automates mixed-precision tuning for HPC applications by leveraging runtime profiling, instruction-level granularity, and temporal locality to learn numerical behavior and recurring patterns in instruction execution. FPLearner (10.1145/3597503.3623338, ) leverages a graph representation learning technique to predict the performance and accuracy of a mixed-precision program, which can then be used to accelerate dynamic precision tuning by reducing the number of program runs.
The aforementioned methods target floating-point programs, and the general ability to handle unseen data remains unverified. Our proposed method operates in a different setting. It proceeds by modeling the rewards associated with multiple objectives and user-defined states (characteristics for the method to be defined in mixed-precision) and learning via several rounds of training on a specific, limited dataset. This approach can be generalized to a variety of unseen testing data. To the best of our knowledge, we are the first to present a work on precision autotuning for numerical algorithms that has verified generalization performance. In the following section, we will detail our framework and methodology for using RL to intelligently select precision configurations.
3 A General RL Framework for Precision Selection
[…]
Figure 1: Schematic of RL Framework: Phase I (top) depicts the iterative training loop where the agent updates the Q-table. Phase II (bottom) shows the inference workflow using the trained RL model.
We formalize the precision selection problem for a general algorithm $\mathcal{M}$ for solving linear systems as a contextual bandit, where an agent dynamically selects precision configurations for computational steps based on features of the problem instance. This framework is summarized in Figure 1 and can easily performed in an online learning routine. It intelligently optimizes the trade-off between numerical accuracy and computational efficiency, leveraging RL to adapt to diverse problem characteristics.
3.1 Contextual Bandit Formulation
Consider a numerical algorithm or method $\mathcal{M}$ designed to solve a problem instance $p\in\mathcal{P}$, where $\mathcal{P}$ denotes the set of problem instances (e.g., linear systems $Ax=b$). The method comprises $k$ computational steps, each requiring a precision choice from a finite set $\mathcal{A}_{i}=\{u_{i,1},u_{i,2},\ldots,u_{i,m}\}$, for $i=1,\ldots,k$, where $m=|\mathcal{A}_{i}|$ is the number of available precisions. The action space is the Cartesian product:
$\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{k},$ (1)
with cardinality $|\mathcal{A}|=m^{k}$. An action is a tuple $\mathbf{a}=(a_{1},a_{2},\ldots,a_{k})\in\mathcal{A}$, where $a_{i}\in\mathcal{A}_{i}$ specifies the precision for the $i$-th computational step. To reduce the computational complexity and simplify the probkem, we assume $|\mathcal{A}_{1}|=|\mathcal{A}_{2}|=\ldots=|\mathcal{A}_{k}|$.
The contextual bandit is defined by:
  • Context ($s\in\mathcal{S}$): A feature vector describing the problem instance $p$, i.e., $s=[\phi_{1}(p),\phi_{2}(p),\ldots,\phi_{d}(p)]\in\mathbb{R}^{d}$, where $\phi_{j}:\mathcal{P}\to\mathbb{R}$ are features capturing properties such as condition number, sparsity, matrix norm, or matrix size.
  • Action ($\mathbf{a}\in\mathcal{A}$): A precision configuration $\mathbf{a}=(a_{1},\ldots,a_{k})$, where $a_{i}$ determines the precision for the $i$-th step.
  • Reward ($r(s,\mathbf{a})$): A scalar evaluating the performance of $\mathcal{M}$ on instance $p$ with action $\mathbf{a}$, defined as a weighted sum of $J$ objectives:
    $r(s,\mathbf{a})=\sum_{j=1}^{J}w_{j}f_{j}(s,\mathbf{a}),$ (2)
    where $f_{j}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are objective functions (e.g., accuracy, computational cost, convergence indicators), and $w_{j}>0$ are weights balancing these objectives.
The goal is to learn a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$ that maximizes the expected reward over a distribution of problem instances $\mathcal{D}$:
$J(\pi)=\mathbb{E}_{s\sim\mathcal{D},\mathbf{a}\sim\pi(\cdot|s)}[r(s,\mathbf{a})].$
3.2 Learning with Discretized State Space
To learn the policy $\pi$ for determining precision choice, we employ a contextual bandit learning paradigm, discretizing the continuous state space $\mathcal{S}$ into a finite set to manage computational complexity. For each feature $\phi_{j}$ ($j=1,\ldots,d$), we define bins $\Phi_{j}=\{b_{j,1},b_{j,2},\ldots,b_{j,n_{j}}\}$, where $n_{j}$ is the number of bins (e.g., logarithmic bins for condition number and norm, linear bins for sparsity and diagonal dominance). The discretized state space is:
$\mathcal{S}_{d}=\Phi_{1}\times\Phi_{2}\times\cdots\times\Phi_{d},$ (3)
with cardinality $|\mathcal{S}_{d}|=\prod_{j=1}^{d}n_{j}$. A state $s_{d}\in\mathcal{S}_{d}$ is obtained by mapping the continuous context $s=[\phi_{1},\ldots,\phi_{d}]$ to a discrete index via binning:
$s_{d}=\text{discretize}(s)=\left(\text{bin}(\phi_{1}(s)),\ldots,\text{bin}(\phi_{d}(s)))\right),$ (4)
where $\text{bin}(\cdot)$ assigns each feature to the nearest bin in $\Phi_{j}$, with clipping to ensure indices remain within bounds.
To handle the large action space $|\mathcal{A}|=m^{k}$, we decompose the action selection by maintaining $k$ separate Q-tables, $Q_{i}:\mathcal{S}_{d}\times\mathcal{A}_{i}\to\mathbb{R}$, for $i=1,\ldots,k$, where $Q_{i}(s_{d},a_{i})$ estimates the expected reward for selecting precision $a_{i}\in\mathcal{A}_{i}$ for the $i$-th step in state $s_{d}$. This assumes independence between precision choices, reducing the Q-table size from $|\mathcal{S}_{d}|\cdot m^{k}$ to $k\cdot|\mathcal{S}_{d}|\cdot m$.
During training, the policy selects actions using an epsilon-greedy strategy to balance exploration and exploitation. For each episode, given a discretized state $s_{d}\in\mathcal{S}$, the agent selects an action $\mathbf{a}\in\mathcal{A}$, where $\mathbf{a}=(u_{1},u_{2},\ldots,u_{k})$ represents a tuple of precisions. The set $\mathcal{A}$ contains (valid) precision combinations. The probability of selecting an action $\mathbf{a}$ is given by:
$\pi(\mathbf{a}|s_{d})=\begin{cases}1-\epsilon+\epsilon/|\mathcal{A}|&\text{if }\mathbf{a}=\arg\max_{\mathbf{a}^{\prime}\in\mathcal{A}}Q(s_{d},\mathbf{a}^{\prime}),\\ \epsilon/|\mathcal{A}|&\text{otherwise},\end{cases}$ (5)
where $\epsilon\in[0,1]$ is the exploration parameter, initialized to $\epsilon_{0}=1.0$ and decayed as $\epsilon\leftarrow\max(\epsilon_{\text{min}},1.0-\text{episode}/\text{total\_episodes})$ with a minimum exploration rate $\epsilon_{\text{min}}$.
After selecting an action $\mathbf{a}$, the environment applies the mixed-precision algorithm to the linear system $A\mathbf{x}=\mathbf{b}$, yielding a reward $R(s_{d},\mathbf{a})$ based on accuracy, precision usage, convergence, and iteration count. The Q-table is updated with a one-step:
$Q(s_{d},\mathbf{a})\leftarrow Q(s_{d},\mathbf{a})+\alpha\Big(R(s_{d},\mathbf{a})-Q(s_{d},\mathbf{a})\Big),$ (6)
where $\alpha\in(0,1]$ is the learning rate. The Q-values are not necessarily clipped in the implementation, relying on the reward function to manage numerical stability.
During inference, the agent selects the action that maximizes the Q-value for the given state:
$\mathbf{a}^{*}=\arg\max_{\mathbf{a}^{\prime}\in\mathcal{A}}Q(s_{d},\mathbf{a}^{\prime}).$ (7)
Algorithm 1 RL Framework for Precision Selection in Numerical Algorithms
1:Input: Training instances $\mathcal{P}=\{p_{1},\ldots,p_{N}\}$, numerical method $\mathcal{M}$, precision sets $\{\mathcal{A}_{i}\}_{i=1}^{k}$, reward function $r$, episodes $T$, learning rate $\alpha$, discount factor $\gamma$, exploration epsilon $\epsilon$, bin sets $\{\Phi_{j}\}_{j=1}^{d}$
2:Initialize Q-table $Q(s_{d},i,a_{i})$ for all $s_{d}\in\mathcal{S}_{d}$, $i=1,\ldots,k$, $a_{i}\in\mathcal{A}_{i}$
3:Initialize environment $\mathcal{E}$ with instances $\{p_{i}\}$
4:for episode $t=1$ to $T$ do
5:  for each instance $p_{i}$, $i=1,\ldots,N$ do
6:   Compute state in terms of data features $s_{i}\leftarrow[\phi_{1}(p_{i}),\ldots,\phi_{d}(p_{i})]$ (e.g., log condition number, log norm)
7:   Discretize state $s_{d}\leftarrow\text{discretize}(s_{i},\{\Phi_{j}\})$
8:   Choose action $a\in\mathcal{A}$ uniformly with probability $\epsilon$ otherwise use $a\leftarrow\arg\max_{a^{\prime}\in\mathcal{A}}Q(s_{d},a^{\prime})$
[…]
10:   Run $\mathcal{M}$ with precision $a$ in $\mathcal{E}$, observe reward $r(s_{i},a)$, next state $s_{i}^{\prime}$, termination status
11:   Discretize next state $s_{d}^{\prime}\leftarrow\text{discretize}(s_{i}^{\prime},\{\Phi_{j}\})$
12:   Update Q-table:
$Q(s_{d},\mathbf{a})\leftarrow=Q(s_{d},\mathbf{a})+\alpha\Big(R(s_{d},\mathbf{a})-Q(s_{d},\mathbf{a})\Big),$
13:   if termination status is true then
14:     Reset $\mathcal{E}$ to next instance
15:   end if
16:  end for
17:  Update epsilon $\epsilon$
18:end for
19:Inference: For new instance $p$, compute $s=[\phi_{1}(p),\ldots,\phi_{d}(p)]$, discretize $s_{d}=\text{discretize}(s,\{\Phi_{j}\})$, select $a_{i}=\arg\max_{a\in\mathcal{A}_{i}}Q(s_{d},i,a)$, apply $\mathcal{M}$ with precision $\mathbf{a}$
Our RL learning framework for precision selection can be extended to general algorithms by properly remodeling the state and action space. The pseudocode the linear solve is detailed in Algorithm 1. The framework aims to optimize precision usage for the $k$ computational steps of a mixed-precision algorithm applied to linear systems $p=(A,b)\in\mathcal{P}$. It iterates over a training set of $N$ problem instances for $T$ episodes, selecting actions according to the estimated Q-values. For each problem instance, it begins by computing a context $s\in\mathbb{R}^{d+n}$ from matrix properties, discretizing it into $s_{d}\in\mathcal{S}_{d}$ using bins $\Phi_{j}$. Actions $\mathbf{a}=(a_{1},\ldots,a_{k})\in\mathcal{A}$ are selected via a softmax policy $\pi(a_{i}|s_{d})\propto\exp(\text{clip}(Q_{i}(s_{d},a_{i}),-Q_{\text{max}},Q_{\text{max}})/\tau)$, where $k$ Q-tables $Q_{i}$ guide precision choices for each step. The solver $\mathcal{M}$ is applied with $\mathbf{a}$, yielding a reward $r(s,\mathbf{a})$ that balances accuracy, cost, convergence, and penalties, clipped to $r_{\text{max}},r_{\text{max}}]$. The Q-tables are updated as $\text{clip}(Q_{i}(s_{d},a_{i})+\alpha(\text{clip}(r(s,\mathbf{a}),-r_{\text{max}},r_{\text{max}})+\gamma\max_{a^{\prime}\in\mathcal{A}_{i}}Q_{i}(s_{d}^{\prime},a^{\prime})-Q_{i}(s_{d},a_{i})),-Q_{\text{max}},Q_{\text{max}})$, iterating over $T$ episodes and $N$ instances with decaying temperature $\tau\leftarrow\max(\tau_{\text{min}},\tau\cdot\gamma_{\tau})$.
In the inference phase, for a new instance $p$, the context $s$ is computed and discretized to $s_{d}$, and precisions are chosen greedily: $a_{i}=\arg\max_{a\in\mathcal{A}_{i}}Q_{i}(s_{d},a)$. The use of separate Q-tables reduces complexity, while clipping ensures numerical stability for ill-conditioned systems. The softmax policy and feature-based context enable the framework to adaptively optimize precision, leveraging matrix properties to achieve efficient and accurate solutions, as validated in numerical experiments in Section 5.
Our RL learning framework optimizes both arithmetic efficiency (by favoring lower-precision steps) and solver performance (e.g., convergence and accuracy). In Section  4, we will apply this framework to GMRES-IR, and detail how we model the reward function based on the earnings and cost for the steps following low precision.
3.3 Action Space Reduction
If the set of precisions is too large, then the size of the action space will exponentially increase, which significantly increases the training load. Let $\mathcal{A}_{1}=\{u_{1},u_{2},\ldots,u_{m}\}$ be the set of precisions ordered by increasing significand bits. For the solver $\mathcal{M}$ and $k$ computational steps for which precisions must be selected, the action space $\mathcal{A}=\mathcal{A}_{1}^{k}$ has cardinality $m^{k}$, where $m$ is the number of precisions available to be chosen. To manage this exponential growth, we prune the action space by imposing a precision order constraint and limiting to the top-$k_{\text{top}}$ combinations.
Here, we write $u_{j}<u_{j^{\prime}}$ if $u_{j}$ has fewer significand bits than $u_{j^{\prime}}$. We enforce that the selected precisions for the $k$ computational steps satisfy
$u^{\prime}_{1}\leq u^{\prime}_{2}\leq u^{\prime}_{3}\leq\ldots\leq u^{\prime}_{k},\qquad u^{\prime}_{j}\in\mathcal{A}_{1}$ (8)
This can naturally lead to valid precision combinations for some algorithms that is meaningful, as discussed in vi22 . This idea is similar to modified delta-debugging in 6877460 , which uses a greedy search approach instead of seeking a global minimum; it might miss potential reduced-precision configurations that suit the algorithm. The principle also naturally enables early stopping. For example, if precision $u^{\prime}_{i}$ fails, then it is not necessary to try a lower precision in step $i$ or to try lower precision for steps $j>i$ if it is already employed. Another benefit is that it allows users to define the precision selection for the $k$ computational steps in terms of their importance (the most important one is assigned to the last step $k$). One benefit of using this strategy is to reduce the action space to
$|\mathcal{A}_{\text{reduced}}|=\binom{m+k-1}{k}=\frac{m!}{k!(m-k)!}.$ (9)
In the following experiments, we prune the action space from 256 to 35, a reduction of approximately 86%.
The exploration strategy is epsilon-greedy with linear decay. For episode $t$, the probability of selecting a random action from $\mathcal{A}_{\text{reduced}}$ is
$\epsilon_{t}=\max(\epsilon_{\text{min}},1.0-t/T).$ (10)
This ensures exploration of lower-precision configurations early in training, shifting to exploitation of the best-known actions as $\epsilon_{t}$ decreases. The reduced action space $\mathcal{A}_{\text{reduced}}$ focuses on numerically stable configurations, accelerating convergence of the RL agent for the solver.
4 Case Study: GMRES-Based Iterative Refinement
4.1 Mixed-Precision GMRES-based Iterative Refinement
Iterative numerical methods, such as those for solving linear systems $Ax=b$, rely on a careful choice of arithmetic precisions to balance computational efficiency and numerical accuracy. The choice of precision for each computational step (e.g., working precision, residual computation) significantly impacts performance, yet optimal selection is challenging due to the complex dependence on system properties, such as condition number and sparsity. Traditional approaches often employ fixed or heuristic precision settings, which may lead to suboptimal performance across diverse problem instances. RL provides a data-driven approach to adaptively select precisions based on system characteristics, optimizing multiple objectives, such as residual norm and iteration count.
We instantiate the general RL framework for the GMRES-IR method (doi:10.1137/17M1140819, ), selecting precisions for four computational steps to solve linear systems $Ax=b$, reconciling the competing demands of numerical accuracy and computational efficiency. Consider a linear system $Ax=b$, where $A\in\mathbb{R}^{n\times n}$ is non-singular, and $b\in\mathbb{R}^{n}$. GMRES-IR iteratively refines an initial solution $x_{0}$ as:
$x_{i+1}=x_{i}+z_{i},\quad\text{where}\quad Az_{i}=r_{i},$
and $r_{i}=b-Ax_{i}$. The system $Az_{i}=r_{i}$ can be solved via preconditioned GMRES, i.e., we solve $M^{-1}Az_{i}=M^{-1}r_{i}$, where $M\approx A$ is a preconditioner. In this study, we use the preconditioner $M=LU$ (typically performed by backward and forward substitution) via LU factorization. The algorithm is presented in Algorithm 2. We refer to the loop in line 3 as outer iterations and the iterations within the GMRES solve in line 5 as inner iterations.
Our mixed-precision algorithms focus on four computational steps, each requiring a precision choice:
  1. 1.
    LU Factorization and Initial Solution: Compute $M=LU\approx A$ and solve $Mx_{0}=b$ in precision $u_{f}$,
  2. 2.
    Residual Computation: Compute the residual $r_{i}=b-Ax_{i}$ in precision $u_{r}$.
  3. 3.
    Working Precision for GMRES: Solve $M^{-1}Az_{i}=M^{-1}r_{i}$ using GMRES in precision $u_{g}$.
  4. 4.
    Solution Update: Compute $x_{i+1}=x_{i}+z_{i}$ in precision $u$.
Algorithm 2 Mixed-precision GMRES based iterative refinement (GMRES-IR)
1:an $n\times n$ matrix $A$ and a right-hand side $b$; $u_{f}$ is precision for factorization;
2:Approximate solution $\hat{x}$ to $A\hat{x}=b$.
3:Compute the LU factorization $A=LU$. $u_{f}$
4:Initialize $x_{0}$ (to, e.g., $x_{0}=U^{-1}L^{-1}(b)$). $u_{f}$
5:for $i=0,\dots$, converged do
6:  Compute $r_{i}=b-Ax_{i}$. $u_{r}$
7:  Solve $U^{-1}L^{-1}Az_{i}=U^{-1}L^{-1}r_{i}$ using GMRES. $u_{g}$
8:  Compute $x_{i+1}=x_{i}+z_{i}$. $u$
9:end for
We note that it is also possible to use mixed precision within the GMRES method itself; see, e.g., vi22 . For simplicity, we restrict ourselves here to GMRES in uniform precision. The goal is to use RL to select an action $\mathbf{a}=(u_{f},u_{r},u_{g},u)$ to minimize the final residual norm $\|b-Ax_{\text{solve}}\|_{2}$ and the total number of iterations $T$, subject to a trade-off between $T_{\text{max}}$ and the convergence tolerance $\tau$.
Inspired by stopping criteria from 10.1145/1141885.1141894 , we consider the algorithm converged if:
  1. 1.
    Convergence: The solver converges if the relative update norm is sufficiently small:
    $\frac{\|z_{i}\|_{\infty}}{\|x_{i}\|_{\infty}}\leq u_{\text{work}},$ (11)
    where $u_{\text{work}}$ is the unit roundoff for working precision, ensuring the update is on the order of the highest precision’s roundoff error.
  2. 2.
    Stagnation: The solver stops if the update norms indicate insufficient progress:
    $\frac{\|z_{i}\|_{\infty}}{\|z_{i-1}\|_{\infty}}\geq\tau,$ (12)
    where $\tau>0$ is the stagnation tolerance, indicating that the updates are not changing significantly.
  3. 3.
    Maximum Iterations: The solver stops if the number of iterations reaches a predefined limit:
    $i\geq i_{\text{max}},$ (13)
    where $i_{\text{max}}$ is the maximum number of iterations.
To evaluate the accuracy and stability of the linear solve, we define the normwise relative forward error, and normwise relative backward error:
$\texttt{ferr}=\frac{\|x_{\text{solve}}-x_{\text{true}}\|_{\infty}}{\|x_{\text{true}}\|_{\infty}},\quad\texttt{nbe}=\frac{\|b-Ax_{\text{solve}}\|_{\infty}}{\|A\|_{\infty}\|x_{\text{solve}}\|_{\infty}+\|b\|_{\infty}}.$ (14)
4.2 Contextual Bandit Formulation for Precision selection
The contextual bandit for the iterative refinement solver with preconditioned GMRES is defined by specifying the context, action, and reward components, tailored to optimize precision selection for solving linear systems $Ax=b$. The context, denoted $s\in\mathbb{R}^{d}$, where $d=2$, captures properties of the linear system and is given by:
$s=[\log_{10}(\max(\kappa(A),\delta_{c})),\log_{10}(\max(\|A\|_{\infty},\delta_{n}))],$ (15)
where $\kappa(A)$ is the condition number of $A$ that can be approximated via an efficient algorithm (e.g., Hager-Higham Hager:1984:CE ; Higham:1987:SCN ), $\delta_{c}>0$ ensures numerical stability in logarithmic scaling, $\|A\|_{\infty}=\max_{i}\sum_{j}|a_{ij}|$ is the infinity norm, and $\delta_{n}>0$ prevents numerical issues. These features characterize the numerical stability and scaling of the linear system and are fast to compute.
To learn the policy $\pi$, we employ a contextual bandit, discretizing the continuous state space $\mathcal{S}\subset\mathbb{R}^{d}$ into a finite set to manage computational complexity. For each feature $\phi_{j}$, $j=1,\ldots,d$, we define bins $\Phi_{j}$. The continuous state space $\mathcal{S}\subset\mathbb{R}^{d}$ is discretized into a finite set $\mathcal{S}_{d}$, as in (3). For features $\phi_{1}=\log_{10}(\max(\kappa(A),\delta_{c}))$ and $\phi_{2}=\log_{10}(\max(\|A\|_{\infty},\delta_{n}))$, we define bins:
  • $\Phi_{1}=\log([\phi_{1,\text{min}},\phi_{1,\text{max}}])$, divided into $n_{1}$ bins for the condition number (e.g., use approximate 1-norm).
  • $\Phi_{2}=\log([\phi_{2,\text{min}},\phi_{2,\text{max}}])$, divided into $n_{2}$ bins for the matrix norm.
The discretized state space is $\mathcal{S}_{d}=\Phi_{1}\times\Phi_{2}$, with cardinality $|\mathcal{S}_{d}|=n_{1}\cdot n_{2}$. A state $s_{d}\in\mathcal{S}_{d}$ is computed as:
$s_{d}=\text{discretize}(s)=\left(\text{bin}(\phi_{1}(s)),\text{bin}(\phi_{2}(s))\right),$ (16)
where $\text{bin}(\phi_{j}(s))$ assigns $\phi_{j}(s)$ to the nearest bin in $\Phi_{j}$, clipped to $0,|\Phi_{j}|-1]$. The state index is:
$s_{d}=\text{bin}(\phi_{1})\cdot n_{2}+\text{bin}(\phi_{2}).$ (17)
The action $\mathbf{a}\in\mathcal{A}$ is a tuple $\mathbf{a}=(u_{f},u,u_{g},u_{r}),\quad u_{i}\in\mathcal{A}_{1}$ where $\mathcal{A}_{1}=\{u_{1},\ldots,u_{m}\}$ is the set of precisions, and $m=|\mathcal{A}_{1}|$. Following (9), the action space $\mathcal{A}=\mathcal{A}_{1}^{4}$ is constrained by:
$u_{f}\leq u\leq u_{g}\leq u_{r},$ (18)
where $\leq$ denotes ordering by increasing significand bits.
The reward $r(s,\mathbf{a})$ has the form
$R(s_{d},\mathbf{a})=w_{2}f_{\text{precision}}(s_{d},\mathbf{a})+w_{1}f_{\text{accuracy}}(s_{d},\mathbf{a})-f_{\text{penalty}},$ (19)
where the weights $w_{1},w_{2}>0$ balance accuracy and cost, and the individual components are defined as follows:
  • Precision term: $f_{\text{precision}}(s_{d},\mathbf{a})$ encourages the use of lower-precision formats based on their significand bits. For each precision $p\in\{u_{f},u,u_{g},u_{r}\}$, the cost is weighted by the ratio of significand bits relative to FP64:
    $f_{\text{precision}}(s_{d},\mathbf{a})=\sum_{p\in\{u_{f},u,u_{g},u_{r}\}}\frac{t_{\text{FP64}}}{t_{p}(1+\log_{10}(\max(\kappa,1)))},$ (20)
    where $t_{u}$ is the number of significand bits for precision $u$, and $\kappa$ is the condition number of the matrix $\mathbf{A}$.
  • Accuracy term: To define the accuracy term, we further refine the error terms based on (14):
    $\text{err}_{\text{distance}}=\frac{\|x_{\text{solve}}-x_{\text{true}}\|_{\infty}}{\|x_{\text{true}}\|_{\infty}},\quad\text{err}_{\text{normalized}}=\frac{\texttt{ferr}}{\|A\|_{\infty}\|x_{\text{true}}\|_{\infty}+\|b\|_{\infty}},$ (21)
    where $x_{\text{solve}}$ is the solution obtained by the reinforcement learning policy, $x_{\text{true}}$ is the ground-truth solution or high-precision reference solution (here we use the one computed in FP64), $\|A\|_{\infty}$ is the matrix infinity norm, and small constants can be added in denominators to prevent division by zero.
    $f_{\text{accuracy}}(s_{d},\mathbf{a})$ penalizes the error in the solution relative to the true solution and the normalized residual. It is computed as:
    $f_{\text{accuracy}}(s_{d},\mathbf{a})=-\log_{10}(\max(\text{err}_{\text{distance}},\mathbf{C}_{1}))-\log_{10}(\max(\text{err}_{\text{normalized}},\mathbf{C}_{1})).$ (22)
    If $\text{err}_{\text{distance}}>1.0\text{ or }\text{err}_{\text{normalized}}>1.0$ we set $f_{\text{accuracy}}(s_{d},\mathbf{a})=\mathbf{C}_{2}$. Empirically, we choose $\mathbf{C}_{1}=10^{-10}$ and $\mathbf{C}_{2}=5.0$.
  • Penalty: We use $f_{\text{penalty}}$ term to denote the penalty arising from computational cost, e.g., iterations, or failure steps such as LU factorization or stagnation calculation. In our empirical study, it is sufficient to use the converged iterations to model the penalty term; that is, the more iterations we use, the larger the penalty term will be. In this context, we can either use the total number of iterations for iterative refinement or the total (inner-solve) GMRES iterations used for iterative refinement, denoted by $T_{\text{iter}}$. The following definition is used in our context:
    $f_{\text{penalty}}=\log_{2}(\max(T_{\text{iter}},1)),$ (23)
    which can be also used in other linear solve. One can also enforce a weight $w_{3}$ on this term to avoid hiding the effect of other terms during training.
We use a contextual bandit update for a single Q-table $Q:\mathcal{S}_{d}\times\mathcal{A}\to\mathbb{R}$. Actions are selected using an epsilon-greedy strategy with
$\epsilon_{t}=\max(\epsilon_{\text{min}},1.0-t/T),$ (24)
as in Section 3.2. The Q-table is updated as
$Q(s_{d},\mathbf{a})\leftarrow Q(s_{d},\mathbf{a})+\alpha\Big(R(s_{d},\mathbf{a})-Q(s_{d},\mathbf{a})\Big),\qquad\mathbf{a}\in\mathcal{A}_{\text{reduced}}.$ (25)
The precision selection RL framework specialized for GMRES-IR is presented in Algorithm 3. It applies RL to a training set of $N$ linear systems, selecting precisions via an epsilon-greedy strategy and updating the Q-table based on solver performance. The GMRES method is adapted to handle mixed-precision computations, with the preconditioner applied in precision $u_{g}$.
Algorithm 3 Contextual Bandit Learning for GMRES-IR Precision Selection
1:Input: Training systems $\{(A_{j},b_{j})\}_{j=1}^{N}$, GMRES-IR method $\mathcal{M}$, precision set $\mathcal{A}_{1}=\{u_{1},\ldots,u_{m}\}$, reward function $r$, episodes $T$, tolerance $\tau$, max iterations $T_{\text{max}}$, learning rate $\alpha$, discount factor $\gamma$, minimum exploration probability $\epsilon_{\text{min}}$, bins $\{\Phi_{1},\Phi_{2}\}$, top-$k$ parameter $k_{\text{top}}$
2:Compute $\mathcal{A}_{\text{reduced}}$ in terms of $\mathcal{A}_{1}$
3:Initialize Q-table $Q(s_{d},\mathbf{a})$ for $s_{d}\in\mathcal{S}_{d}$, $\mathbf{a}\in\mathcal{A}_{\text{reduced}}$
4:Initialize environment $\mathcal{E}$ with systems $\{(A_{j},b_{j})\}$
5:for episode $t=1$ to $T$ do
6:  for system $(A_{j},b_{j})$, $j=1,\ldots,N$ do
7:   Compute context $s_{j}=[\log_{10}(\max(\kappa(A_{j}),\delta_{c})),\log_{10}(\max(\|A_{j}\|_{\infty},\delta_{n}))]$
8:   Discretize state $s_{d}=\text{discretize}(s_{j},\{\Phi_{1},\Phi_{2}\})$
9:   Compute $\epsilon_{t}=\max(\epsilon_{\text{min}},1.0-t/T)$
10:   Select action $\mathbf{a}=(u_{f},u,u_{g},u_{r})\in\mathcal{A}_{\text{reduced}}$ via:
$\mathbf{a}=\begin{cases}\text{random choice from }\mathcal{A}_{\text{reduced}},&\text{if }\text{with }\epsilon_{t}\text{ probability},\\ \arg\max_{\mathbf{a}^{\prime}\in\mathcal{A}_{\text{reduced}}}Q(s_{d},\mathbf{a}^{\prime}),&\text{otherwise}.\end{cases}$
11:   Solve $A_{j}x=b_{j}$ with $\mathcal{M}$ using $\mathbf{a}$:
12:    Compute $M=LU\approx A_{j}$ in precision $u_{f}$
13:    Solve $Mx_{0}=b_{j}$ in precision $u_{f}$
14:    for $l=0$ to $T_{\text{max}}-1$:
15:     Compute $r_{l}=b_{j}-A_{j}x_{l}$ in precision $u_{r}$
16:     Solve $M^{-1}A_{j}z_{l}=M^{-1}r_{l}$ using GMRES with in precision $u_{g}$
17:     Update $x_{l+1}=x_{l}+z_{l}$ in precision $u$
18:     if stopping criteria is met or failure occurs: break
19:    Compute reward $r(s_{j},\mathbf{a})$ per (19)
20:   Compute next system index $j^{\prime}=(j\bmod N)+1$, context $s^{\prime}$, and discretized state $s_{d}^{\prime}$
21:   Update Q-table:
$Q(s_{d},\mathbf{a})\leftarrow=Q(s_{d},\mathbf{a})+\alpha\Big(R(s_{d},\mathbf{a})-Q(s_{d},\mathbf{a})\Big)$
22:  end for
23:end for
24:Inference: For new system $(A,b)$, compute $s=[\log_{10}(\max(\kappa(A),\delta_{c})),\log_{10}(\max(\|A\|_{\infty},\delta_{n}))]$, discretize $s_{d}$, select $\mathbf{a}=\arg\max_{\mathbf{a}\in\mathcal{A}_{\text{reduced}}}Q(s_{d},\mathbf{a})$, solve with $\mathcal{M}$ using $\mathbf{a}$
5 Experiments
We conducted experiments to evaluate the proposed RL framework for precision selection in GMRES-IR for solving linear systems $p=(A,b)\in\mathcal{P}$, where $A\in\mathbb{R}^{n\times n}$ is a non-singular matrix and $b\in\mathbb{R}^{n}$. The systems are derived from random generated matrices. Our code is simulated in Python and uses Pychop (carson2025, ) for precision emulation.
We generated problems $Ax=b$ using two settings, each designed to evaluate the RL-driven precision selection for the GMRES-IR solver under different numerical conditions versus a full double precision routine (in the following, we use FP64 and double precision interchangeably). The agent with learning rate $\alpha$ of $0.5$ is trained within a custom Gym environment (brockman2016openai, ) with 100 episodes, which simulates the process of solving a linear system using the GMRES-IR solver. To reduce the action space and accelerate computation, we adopted the reduction strategy outlined in Section 3.3 and further pruned it by selecting one-fourth of the valid precision combinations from the episode number combinations.
In the following experiments, we evaluate the resulting relative forward errors (as defined in (14)), the number of iterative refinement iterations required for convergence, and the total number of inner GMRES iterations required for convergence. All the computed results for the relative forward and backward errors (abbreviated as error in the tables and figures for simplicity in the following) are preserved to 2 significant digits. For a fair comparison, we set the same convergence tolerance $\tau$ for both RL baseline and the reference baseline.
5.1 Problem Setup
To ensure sufficient diversity in the data, we synthetically generated the linear systems, comprising $N_{\text{train}}$ training and $N_{\text{test}}$ testing systems $\{\mathcal{P}_{i}=(A_{i},b_{i})\}_{i=1}^{N_{\text{train}}+N_{\text{test}}}$. The data generation iterated until $N_{\text{train}}+N_{\text{test}}=200$ systems were obtained. Specifically, we set $N_{\text{train}}=100$ training and $N_{\text{test}}=100$ testing systems. The generation process incorporates randomization for matrix sizes (between 100 and 500) and condition number ($10^{1},10^{9}$) to ensure diversity in matrix structures, enforcing variability in numerical stability. This diversity is critical for evaluating our RL-based precision selection strategy. We use precision set $\mathcal{U}=\{\text{BF16},\text{TF32},\text{FP32},\text{FP64}\}$ and the discretized state as in (3), based on two features: $\phi_{1}=\log_{10}(\max(\kappa(A),\delta_{c}))$ (condition number) and $\phi_{2}=\log_{10}(\max(\|A\|_{\infty},\delta_{n}))$ (matrix norm), where $\delta_{c},\delta_{n}>0$ ensure numerical stability. These features $\Phi_{1}$ and $\Phi_{2}$ are separately binned into 10 bins ($n_{1}=n_{2}=10$), respectively and their minimum and maximum value in terms of the training set. The discretized state space is $\mathcal{S}_{d}=\Phi_{1}\times\Phi_{2}$, with cardinality $|\mathcal{S}_{d}|=n_{1}\cdot n_{2}$, and states are indexed as in (3).
To evaluate the performance of the RL-based solver $\mathcal{M}$, we define the success rate across a test set of linear systems $\{\mathcal{P}_{1},\ldots,\mathcal{P}_{M}\}\subset\mathcal{S}$. For each system $\mathcal{P}=(A,b)$, the solver produces a solution $x_{\text{rl}}$ using the learned policy $\pi$.
The performance is assessed using the following error metrics maximum error ferr and nbe, defined by:
$\epsilon_{\text{max}}(\mathcal{P},\mathbf{a})=\max(\texttt{ferr},\texttt{nbe}).$
For each condition number range $\mathcal{R}j$, a threshold is defined:
$\tau_{j}=\tau_{\text{base}}\cdot\text{median}({\kappa(A)\mid\mathcal{P}=(A,b)\in\mathcal{R}j}),$ (26)
where $\tau_{\text{base}}>0$ is the base tolerance. A system $\mathcal{P}\in\mathcal{R}j$ is considered successfully solved if:
$\epsilon{\text{max}}(\mathcal{P},\mathbf{a})<\tau_{j}.$ (27)
To evaluate the proportion of systems that achieve acceptable accuracy, the success rate $\xi_{j}$ for the range $\mathcal{R}_{j}$ is:
$\xi_{j}=\frac{|{\mathcal{P}\in\mathcal{R}j\mid\epsilon{\text{max}}(\mathcal{P},\mathbf{a})<\tau_{j}}|}{|\mathcal{R}_{j}|},$ (28)
where $|\mathcal{R}_{j}|$ is the cardinality $\mathcal{R}_{j}$, and the threshold $\tau_{j}$ is scaled by the condition number to account for problem difficulty, aligning with the error bounds in 10.1145/1141885.1141894 .
To comprehensively assess the performance of our approach and study the hyperparameters, we choose the precisions types of BF16, TF32, FP32, and FP64 for exploration and simulate our models with two weight settings $W_{1}$ and $W_{2}$ for (19); $W_{1}$ indicates $w_{1}=1,w_{2}=0.1$, and $W_{2}$ indicates $w_{1}=w_{2}=1.0$. Compared to $W_{1}$, the setting $W_{2}$ increases the value of second weight parameter, and is expected to enable the model to choose the lower precision more aggressively.
The following simulations evaluate performance using the success rate $\xi$, the normative relative forward error (ferr), the normative relative backward error (nbe), and the iteration count, with comparisons against an FP64 baseline. All the Avg. ferr, Avg. nbe, Avg iter. and Avg GMRES iter. in the tables are separately calculated from the average value of ferr, nbe, converged iterations for iterative refinement, and total GMRES iterations used for inner solve from all samples, respectively. The reward and the average Reward Prediction Error (RPE) per episode during training are shown in the Appendix.
5.2 Simulations on Dense Systems
We generated linear systems with varying condition numbers and matrix sizes for training and testing sets, following MATLAB’s gallery(’randsvd’) function with mode=2. These datasets are created using a singular value decomposition approach. The synthetic sets test solvers with mixed-precision configurations determined by RL. This allows us to verify the reinforcement learning agent’s robustness in selecting precision strategies for systems that exhibit variability. For a matrix $A\in\mathbb{R}^{n\times n}$, we construct:
$A=U\Sigma V^{T},$
where $U,V\in\mathbb{R}^{n\times n}$ are orthogonal matrices obtained via QR decomposition of random matrices with entries drawn from a standard normal distribution, and $\Sigma=\operatorname{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{n})$ is a diagonal matrix with the singular values.
The singular values are:
$\sigma_{k}=\begin{cases}\sigma_{\text{max}},&k=1,2,\ldots,n-1,\\ \sigma_{\text{max}}/\kappa,&k=n,\end{cases}$ (29)
where $\sigma_{\text{max}}>0$ is the largest singular value and $\kappa\geq 1$ is the desired condition number, ensuring $\kappa(A)=\sigma_{1}/\sigma_{n}=\kappa$. The singular values are sorted in descending order ($\sigma_{1}=\sigma_{2}=\cdots=\sigma_{n-1}=\sigma_{\text{max}}\geq\sigma_{n}=\sigma_{\text{max}}/\kappa$), resulting in a dense matrix $A$ with the specified condition number $\kappa(A)=\kappa$. The ground-truth solution $x\in\mathbb{R}^{n}$ is generated with entries independently sampled from a standard normal distribution, and the right-hand side vector $b=Ax$.
Table 2: Average Performance Metrics Across Condition Ranges for Dense Systems.
Method Condition Range $\bm{\xi}$ Avg. ferr Avg. nbe Avg iter. Avg. GMRES iter.
$\tau=10^{-6}$
RL($W_{1}$) Low ($10^{0}$$10^{3}$) 100 % $1.19\text{\times}{10}^{-14}$ $7.90\text{\times}{10}^{-17}$ $2.35$ $2.70$
Medium ($10^{3}$$10^{6}$) 100 % $1.57\text{\times}{10}^{-11}$ $82\text{\times}{10}^{-15} 2.2$62.26 $2.20$
High ($10^{6}$$10^{9}$) 100 % $1.89\text{\times}{10}^{-09}$ $1.63\text{\times}{10}^{-14}$ $2.21$ $2.54$
RL($W_{2}$) Low ($10^{0}$$10^{3}$) 100 % $2.48\text{\times}{10}^{-07}$ $2.19\text{\times}{10}^{-08}$ $2.30$ $8.00$
Medium ($10^{3}$$10^{6}$) 100 % $4.22\text{\times}{10}^{-07}$ $4.95\text{\times}{10}^{-11}$ $2.29$ $3.40$
High ($10^{6}$$10^{9}$) 100 % $1.89\text{\times}{10}^{-09}$ $1.63\text{\times}{10}^{-14}$ $2.21$ $2.54$
[…]
$\tau=10^{-8}$
RL($W_{1}$) Low ($10^{0}$$10^{3}$) 100 % $1.18\text{\times}{10}^{-14}$ $7.40\text{\times}{10}^{-17}$ $2.35$ $4.49$
Medium ($10^{3}$$10^{6}$) 100 % $2.95\text{\times}{10}^{-11}$ $9.60\text{\times}{10}^{-13}$ $2.14$ $4.77$
High ($10^{6}$$10^{9}$) 100 % $2.02\text{\times}{10}^{-09}$ $7.93\text{\times}{10}^{-17}$ $2.11$ $3.79$
RL($W_{2}$) Low ($10^{0}$$10^{3}$) 89.2 % $2.49\text{\times}{10}^{-07}$ $61\text{\times}{10}^{-08} 2.2$22.22 $9.54$
Medium ($10^{3}$$10^{6}$) 100 % $3.49\text{\times}{10}^{-06}$ $6.44\text{\times}{10}^{-11}$ $2.14$ $4.71$
High ($10^{6}$$10^{9}$) 100 % $3.91\text{\times}{10}^{-06}$ $4.28\text{\times}{10}^{-11}$ $2.11$ $3.57$
[…]
Table 3: Average Floating-point Types Selected Across Condition Ranges for Dense Systems.
Weight Setting Condition Range Num BF16 TF32 FP32 FP64
$\tau=10^{-6}$
RL ($W_{1}$) $10^{0}\!-\!10^{1}$ 33 $0.00$ $0.00$ $0.33$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.00$ $0.00$ $0.43$ $3.57$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.00$ $0.86$ $3.14$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.33$ $3.67$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.78$ $3.22$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
RL ($W_{2}$) $10^{0}\!-\!10^{1}$ 33 $0.33$ $0.00$ $0.00$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.19$ $0.76$ $0.10$ $2.95$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.00$ $0.86$ $3.14$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.23$ $0.23$ $3.54$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.33$ $3.67$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.78$ $3.22$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
$\tau=10^{-8}$
RL ($W_{1}$) $10^{0}\!-\!10^{1}$ 33 $0.00$ $0.00$ $0.33$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.00$ $0.00$ $0.43$ $3.57$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.29$ $0.29$ $3.43$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.20$ $3.80$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.11$ $3.89$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
RL ($W_{2}$) $10^{0}\!-\!10^{1}$ 33 $0.33$ $0.00$ $0.00$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.14$ $0.76$ $0.10$ $3.00$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.29$ $0.29$ $3.43$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.33$ $3.67$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.78$ $3.22$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
Table 5.2 reports error metrics and iteration counts for the linear solve, organized by condition number ranges: low ($10^{0}$ to $10^{3}$), medium ($10^{3}$ to $10^{6}$), and high ($10^{6}$ to $10^{9}$). Table 3 details the precision types selected within finer condition number intervals.
(a) $\tau=10^{-6}$
(b) $\tau=10^{-6}$
Figure 2: Comparison of RL and FP64 in terms of errors and iteration counts. Marker colors denote different matrix-size groups, while marker sizes represent the number of test samples in each group.
In terms of convergence, the conservative $W_{1}$ policy achieves universal convergence with $\xi=100\%$ across all settings. By comparison, the aggressive $W_{2}$ policy maintains $100\%$ success at $\tau=10^{-6}$ but shows reduced robustness at the stricter tolerance $\tau=10^{-8}$, with $\xi$ decreasing to $89.2\%$ in the low-$\kappa$ tests (RL fails in a few test samples), likely because the RL had insufficient exposure to training data in this specific regime. With respect to accuracy, $W_{1}$ closely matches the FP64 baseline, producing forward error with the same order of magnitude and comparable backward errors. Precision lower than FP32 is never used in $W_{1}$. In contrast, $W_{2}$ exhibits a clear trade-off: BF16 and TF32 start to be used within low-to-medium $\kappa$ ranges, where smaller condition numbers theoretically permit lower-precision computations with decreased error amplification. In $W_{2}$, it permits the use of low precisions with substantially higher errors (e.g., ferr $\approx 10^{-7}$) to optimize the reward function, as long as the convergence quality is guaranteed. This behavioral distinction manifests in computational effort as well. The average number of outer iterations remains stable ($\approx 2.0$ to $2.35$) for both policies. However, the inner solver cost varies. The $W_{2}$ policy requires slightly more inner GMRES iterations (e.g., $8.00$ compared to $2.70$ for $W_{1}$ in low-$\kappa$ bins). This increase is needed to offset the growing rounding errors introduced by lower-precision formats. The visualization of ferr and total GMRES iterations used for each test sample in $W_{2}$ is shown in Figure 2. The average performance degradation in Table 5.2 from the $W_{2}$ is offset by a few instances that deviate from the identity line due to the aggressive strategy with reduced-precision preference.
The precision selection patterns in Table 3 clarify these performance outcomes. In the low-$\kappa$ regime ($10^{0}$ to $10^{2}$), $W_{1}$ employs a conservative combination of FP64 and FP32, entirely avoiding lower precisions. In contrast, $W_{2}$ leverages BF16 and TF32 (with usage frequencies of approximately $0.33$ and $0.76$, respectively), reflecting a focus on cost-efficiency. As the condition number increases, both strategies converge; beyond $10^{6}$, both policies adopt an FP64-dominant strategy and avoid the use of lower precisions.
Although our RL framework for precision selection tolerates a loss of accuracy, the primary finding is that the RL agent develops an intelligent condition-dependent precision-adaptation strategy. It utilizes lower-precision formats (BF16/TF32) to achieve computational efficiency in well-conditioned systems, particularly under the $W_{2}$ policy, but transitions decisively to high-precision FP64 as the matrix becomes ill-conditioned. This behavior demonstrates that the RL agent internalizes theoretical constraints on backward error growth and dynamically prioritizes numerical stability over computational speed when problem difficulty increases.
5.3 Simulations on Sparse Systems
In this simulation, we emulated sparse linear systems following the experiment described in hausner2024neural . A matrix $A_{0}\in\mathbb{R}^{n\times n}$ is constructed with $\text{nnz}(A_{0})=\lfloor\lambda_{s}n^{2}\rfloor$ non-zero entries, where $\lambda_{s}>0$ is the sparsity parameter. The entries are sampled from a standard normal distribution:
$a_{ij}\sim\mathcal{N}(0,1),\quad(i,j)\in\{(r_{k},c_{k})\}_{k=1}^{\text{nnz}(A_{0})},$
where $r_{k},c_{k}\in\{1,\ldots,n\}$ are randomly chosen indices. The matrix is symmetrized and made positive definite:
$A=A_{0}A_{0}^{T}+\beta I,$
where $\beta>0$ is a diagonal shift and $I$ is the identity matrix. The same as the tests for dense linear systems, the ground-truth solution $x\in\mathbb{R}^{n}$ is generated with entries independently sampled from a standard normal distribution, and the right-hand side vector $b=Ax$. For all problems, matrices have non-zero diagonal entries ($a_{ii}\neq 0$) and are non-singular.
Table 4: Train/Test Metrics Summary
Metric Train (min – max) Test (min – max)
Condition number $9.926\times 10^{7}$$1.559\times 10^{10}$ $1.016\times 10^{8}$$7.863\times 10^{9}$
Sparsity 1.80% – 5.10% 1.80% – 5.00%
Matrix size 103 – 500 100 – 498
Setting $\lambda_{s}=0.01$, the random sparse linear systems in our training set and testing set following the above generation routines are described in Table 4. Similar to the above (under the same evaluation metrics, two tolerance levels, and two reward weight policies), Table 5 showcases the average performance of the RL-guided mixed-precision iterative refinement solver on sparse linear systems. Table 6 details the average selection frequency of four floating-point formats across the four precision-controlled stages, with each row summing to 4.
The evaluation of sparse linear systems presents a distinct contrast to the dense case, primarily driven by the inherent numerical difficulty posed by the very ill-conditioned dataset. As depicted in Table 4, the test matrices exhibit high condition numbers between $1.02{\times}10^{8}$ and $7.86{\times}10^{9}$, which places linear solve in a regime where numerical stability is precarious, leaving limited room for precision reduction. This test effectively demonstrates whether the RL agent can effectively avoid numerical instabilities in its precision selection.
Table 5 summarizes the performance metrics. Under these harsh conditions, the RL agent exhibits a highly conservative behavior. Regardless of the reward weight setting ($W_{1}$ or $W_{2}$) or the tolerance level ($\tau$), the agent achieves a $100\%$ success rate with error metrics (ferr and nbe) that are close to the FP64 baseline. Unlike the dense case, where $W_{2}$ triggered an increase in GMRES iterations to trade off lower precision, here the average iteration counts (both outer and GMRES) for RL policies align closely with the baseline. This indicates that the agent has learned that any reduction in precision, even for the sake of potential reward, would likely lead to stagnation or divergence.
This conservative approach is further quantified in Table 6. The precision distribution indicates a consistent reliance on double precision, with FP64 usage averaging approximately $3.99$ to $4.00$ steps per solve. Even under the aggressive weighting scheme $W_{2}$, which strongly incentivizes low-precision usage, the agent entirely avoids selecting BF16 and TF32. This convergence of $W_{1}$ and $W_{2}$ strategies is a significant finding: it demonstrates that the RL agent correctly identifies the "survival boundary" of the iterative method. When faced with uniformly ill-conditioned sparse systems, the policy effectively overrides the efficiency-driven incentives of $W_{2}$ and prioritizes the fundamental necessity of numerical convergence. Thus, the framework proves robust, defaulting to a high-precision fallback when the problem’s spectral properties render low-precision formats numerically unsafe.
Table 5: Average Performance Metrics for Sparse Systems.
Method $\bm{\xi(\%)}$ Avg. ferr Avg. nbe Avg Iter. Avg. GMRES iter.
$\tau=10^{-6}$
RL ($W_{1}$) $100.00$ $4.91\text{\times}{10}^{-09}$ $5.26\text{\times}{10}^{-17}$ $2.01$ $2.22$
RL ($W_{2}$) $100.00$ $4.91\text{\times}{10}^{-09}$ $5.26\text{\times}{10}^{-17}$ $2.01$ $2.22$
[…]
$\tau=10^{-8}$
RL ($W_{1}$) $100.00$ $4.90\text{\times}{10}^{-09}$ $5.28\text{\times}{10}^{-17}$ $2.00$ $2.10$
RL ($W_{2}$) $100.00$ $4.91\text{\times}{10}^{-09}$ $5.28\text{\times}{10}^{-17}$ $2.01$ $2.36$
[…]
Table 6: Average floating-point precision usage per solve for sparse systems.
Weight Setting BF16 TF32 FP32 FP64
$\tau=10^{-6}$
RL ($W_{1}$) $0.00$ $0.00$ $0.01$ $3.99$
RL ($W_{2}$) $0.00$ $0.00$ $0.01$ $3.99$
$\tau=10^{-8}$
RL ($W_{1}$) $0.00$ $0.00$ $0.00$ $4.00$
RL ($W_{2}$) $0.00$ $0.00$ $0.01$ $3.99$
6 Discussion
Our RL framework for precision selection requires proper state modeling of the data. For instance, we use features such as the approximate condition number and matrix norm in our state modeling for the linear solve problem. Selecting or modeling the states often requires expert knowledge to balance reasonable computational complexity with sufficient information for effective RL training.
However, this limitation, on the other hand, can also be viewed as advantages. In this work, the proposed framework has shown efficient precision selection even when prior knowledge about the required precision for critical computational steps is lacking. Although prior studies have explored mixed-precision usage in iterative refinement methods based on condition numbers and matrix sizes, our framework offers two unique benefits. First, it can quickly determine the required precision for unseen linear systems. Second, it allows us to identify the required precision for selected matrix features in a black-box setting. For example, by selecting certain features in our state space, we can examine whether these features are key factors (salient or not) that determine the reduced mixed precision. When the salient features are unknown or when their influence on the required precision level is unclear, our framework becomes particularly advantageous.
7 Conclusion and Future Work
We propose a reinforcement learning framework for automated precision selection for linear solvers by modeling a contextual bandit problem, with a case study in iterative refinement. This framework enables adaptive precision selection for customized computational steps, thereby optimizing the trade-off between numerical accuracy and computational efficiency. Our empirical results show that RL, leveraging a discretized state space, effectively adapts precision choices to the numerical properties of each linear system and achieves a good compromise between algorithm performance (relative error and converged iterations) and the reduced-precision portion of the linear systems under varying conditions.
In the future, we will explore deep RL for continuous-state spaces, incorporate additional significant matrix features (e.g., fill-in ratio, pivot growth estimates, sparsity, and spectral properties), and examine their contributions to improving the method’s performance by reducing precision. By addressing these, we will be able to extend this framework to a broader range of numerical algorithms to advance the mixed-precision techniques.